# Artificial barriers for stochastic differential equations and for construction of boundary-preserving schemes

### v2

@@ 6. Artificial Barriers Euler–Peano
where the constant is independent of $b$ if $b=\infty$. Therefore, applying ${\mathbb{E}}$ to equation (39) and using the bounds for ${\mathbb{E}}\left[J_{1}\right]$ and ${\mathbb{E}}\left[J_{2}\right]$, we obtain
@@
$\sup_{t\in[0,T]}|\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}(t)-X(t)|\leq\eta_{\epsilon}\Delta t^{1/2-\epsilon}$
almost surely. Moreover, $\eta_{\epsilon}$ is independent of $N\in\mathbb{N}$.
@@ 7. Numerical experiments
We present numerical experiments to numerically confirm the theoretical results obtained in Theorem 25. Recall from Section 4 that $M\in\mathbb{N}$ denotes the number of subintervals $[t_{m},t_{m+1}]$ each of size $\Delta t=T/M$. In this section we strive to verify that the ABEM scheme achieves $L^{p}(\Omega)$-convergence rate of $1/2-$ as stated in Theorem 25, and to this end we couple $N$ and $\Delta t$ as $N=\Delta t^{-d}$ for some $d>0$. Let $\Delta B_{m}=B(t_{m+1})-B(t_{m})$ denote the Brownian motion increment over the interval $[t_{m},t_{m+1}]$.
@@ 7.1. Geometric Brownian motion
with initial value $X(0)=x_{0}\in(0,\infty)$; that is, $f(x)=-x$ and $g(x)=\lambda x$ are linear in the considered SDE in equation (1). We introduce a parameter $\lambda>0$ to better illustrate that the EM, SEM and TE schemes do not preserve the boundaries of the SDE (41). In this case $D=(0,\infty)$ and the SDE in (41) is an example of a one-finite-boundary case. The solution $X(t)$ of (41) is called geometric Brownian motion and is given by
@@ 7.2. Allen–Cahn type SDE
(43) $\mathop{}\!\mathrm{d}X(t)=\left(X(t)-X(t)^{3}\right)\mathop{}\!\mathrm{d}t+\lambda(1-X(t)^{2})\mathop{}\!\mathrm{d}B(t)$
with initial value $X(0)=x_{0}\in(-1,1)$; that is, $f(x)=x-x^{3}$ is cubic and $g(x)=\lambda(1-x^{2})$ is quadratic in the considered SDE in equation (1). For the same reason as in the example of geometric Brownian motion (gBm) in Section 7.1, we use a parameter $\lambda>0$ in (43). In this case $D=(-1,1)$ and the Allen-Cahn type SDE in (43) represents an instance of a two-finite-boundaries case. The considered Allen-Cahn type SDE is motivated, for example, by a spatial finite difference space discretisation of a stochastic Allen-Cahn PDE [2, 6, 12, 27]. In contrast to the gBm case in Section 7.1, there is no closed form solution to (43).
@@
As in Section 7.1, we present in Table 2 the proportions out of $100$ samples of the EM, SEM, TE and ABEM schemes that preserved the domain $D=(-1,1)$ of the SDE in (43). The outcome of the experiment in Table 2 confirms that the ABEM scheme is boundary-preserving and that the EM, SEM and TE schemes are not. We used the parameters $T=1$, $x_{0}$ uniformly distributed on $(-1,1)$, $d=1/2$ and $\Delta t=10^{-2}$ in Table 2.
@@
$1$ $/100$ 100 $/100$ 100 $/100$ 100 $/100$ 100
$1.5$ $99/100$ $99/100$ $99/100$ $/100$ 100
$2$ $59/100$ $62/100$ $65/100$ $/100$ 100
Table 2. Proportion of samples containing only values in $D=(-1,1)$ out of $100$ simulated sample paths for the Euler–Maruyama scheme (EM), the semi-implicit Euler–Maruyama scheme (SEM), the tamed Euler scheme (TE), and Artificial Barriers Euler–Maruyama scheme (ABEM) for the Allen–Cahn type SDE in (43) for different choices of $\lambda>0$ and parameters $T=1$, $x_{0}$ uniformly distributed on $(-1,1)$ for each sample, $d=1/2$ and $\Delta t=10^{-2}$.
@@ 7.3. Hat function SDE
(44) $\mathop{}\!\mathrm{d}X(t)=-H(X(t))\mathop{}\!\mathrm{d}t+\lambda H(X(t))\mathop{}\!\mathrm{d}B(t)$
@@ Acknowledgements
The author would like to thank Akash Sharma for reading and discussing the first version of the paper. The author would also like to thank David Cohen, Erik Jansson and Ioanna Motschan-Armen for their helpful comments to improve the presentation pf parts of of the work. This work is partially supported by the Swedish Research Council (VR) (David Cohen’s project nr. $2018-04443$). The computations were enabled by resources provided by Chalmers e-Commons at Chalmers.

### v3

@@ 6. Artificial Barriers Euler–Peano
where the constant is independent of $b$ if $b=\infty$. Therefore, applying $\mathbb{E}$} to equation (40) and using the above bounds for ${\mathbb{E}}\left[J_{1}\right]$ and ${\mathbb{E}}\left[J_{2}\right]$, we obtain
@@
$\sup_{t\in[0,T]}|\hat{\xi}^{\textrm{ABEP}}_{N,\Delta t}(t)-X(t)|\leq\eta_{\epsilon}\Delta t^{1/2-\epsilon},$
almost surely. Moreover, $\eta_{\epsilon}$ is independent of $N\in\mathbb{N}$ and $\Delta t$.
@@ 7. Numerical experiments
We present numerical experiments to numerically confirm the theoretical results obtained in Theorem 25. Recall from Section 4 that $M\in\mathbb{N}$ denotes the number of subintervals $t_{m},t_{m+1}]$ each of size $\Delta t=T/M$. In this section we strive to verify that the ABEM scheme achieves $L^{p}(\Omega)$-convergence rate of $1/2$- as stated in Theorem 25, and to this end we couple $N$ and $\Delta t$ as $N=\Delta t^{-d}$ for some $d>0$. Let $\Delta B_{m}=B(t_{m+1})-B(t_{m})$ denote the Brownian motion increment over the interval $t_{m},t_{m+1}]$.
@@ 7.1. Geometric Brownian motion
with initial value $X(0)=x_{0}\in(0,\infty)$; that is, f(x)=xf(x)=-x and $g(x)=\lambda x$ are linear in the considered SDE in equation (1). We introduce a parameter $\lambda>0$ to better illustrate that the EM, SEM and TE schemes do not preserve the boundaries of the SDE (42). In this case $D=(0,\infty)$ and the SDE in (42) is an example of a one-finite-boundary case. The solution $X(t)$ of (42) is called geometric Brownian motion and is given by
@@ 7.2. Allen–Cahn type SDE
(44) $\mathop{}\!\mathrm{d}X(t)=\left(X(t)-X(t)^{3}\right)\mathop{}\!\mathrm{d}t+\lambda(1-X(t)^{2})\mathop{}\!\mathrm{d}B(t),\ t\in[0,T]$
with initial value $X(0)=x_{0}\in(-1,1)$; that is, $f(x)=x-x^{3}$ is cubic and $g(x)=\lambda(1-x^{2})$ is quadratic in the considered SDE in equation (1). For the same reason as in the example of geometric Brownian motion (gBm) in Section 7.1, we use a parameter $\lambda>0$ in (44). In this case D=(1,1)D=(-1,1) and the Allen-Cahn type SDE in (44) represents an instance of a two-finite-boundaries case. The considered Allen-Cahn type SDE is motivated, for example, by a spatial finite difference space discretisation of a stochastic Allen-Cahn PDE [2, 8, 16, 34]. In contrast to the gBm case in Section 7.1, there is no closed form solution to (44).
@@
As in Section 7.1, we present in Table 2 the proportions out of $100$ samples of the EM, SEM, TE and ABEM schemes that preserved the domain D=(1,1)D=(-1,1) of the SDE in (44). The outcome of the experiment in Table 2 confirms that the ABEM scheme is boundary-preserving and that the EM, SEM and TE schemes are not. We used the parameters $T=1$, $x_{0}$ uniformly distributed on (1,1)(-1,1), $d=1/2$ and $\Delta t=10^{-2}$ in Table 2.
@@
11 $100/100$ $100/100$ $100/100$ $100/100$
$1.5$ $99/100$ $99/100$ $99/100$ $100/100$
22 $59/100$ $62/100$ $65/100$ $100/100$
Table 2. Proportion of samples containing only values in D=(1,1)D=(-1,1) out of $100$ simulated sample paths for the Euler–Maruyama scheme (EM), the semi-implicit Euler–Maruyama scheme (SEM), the tamed Euler scheme (TE), and Artificial Barriers Euler–Maruyama scheme (ABEM) for the Allen–Cahn type SDE in (44) for different choices of $\lambda>0$ and parameters $T=1$, $x_{0}$ uniformly distributed on (1,1)(-1,1) for each sample, $d=1/2$ and $\Delta t=10^{-2}$.
@@ 7.3. Hat function SDE
(45) $\mathop{}\!\mathrm{d}X(t)=-H(X(t))\mathop{}\!\mathrm{d}t+\lambda H(X(t))\mathop{}\!\mathrm{d}B(t),\ t\in[0,T],$
@@ Acknowledgements
The author would like to thank Akash Sharma for reading and discussing the first version of the paper. The author would also like to thank David Cohen, Erik Jansson and Ioanna Motschan–Armen for their helpful comments to improve the presentation of parts of the work. This work is partially supported by the Swedish Research Council (VR) (David Cohen’s project nr. 2018044432018-04443). The computations were enabled by resources provided by Chalmers e-Commons at Chalmers.